OpenStudy (anonymous):

The length of a rectangle is one less then 2 times the width. The area of the rectangle is 15 square inches. Find the length and width of this rectangle

OpenStudy (anonymous):

So you know the area=length x width and you know the length is one less than two times the width so let the width=w and length =L so L=2w-1 So area = L x W = (2w-1) x w = 15 so 2w^2-w=15 rearrange and solve for w (ie the with) and then substitute in the find the length